Apolonia:

Use properties to rewrite the given equation. Which equations have the same solution as x + + x = – x? Check all that apply. x + = – x 18x + 20 + 30x = 15 – 6x 18x + 20 + x = 15 – 6x 24x + 30x = –5 12x + 30x = –5

Logic007:

http://prntscr.com/jxnt5y Hmm,something is missing It will do good,if u can double check ur question

Apolonia:

Use properties to rewrite the given equation. Which equations have the same solution as 3/5x + 2/3 + x =1/2 –1/5 x? Check all that apply. 8/5x + 2/3 = 1/2 –1/5 x 18x + 20 + 30x = 15 – 6x 18x + 20 + x = 15 – 6x 24x + 30x = –5 12x + 30x = –5

Logic007:

That's better :)

Logic007:

\(\frac{3}{5}x+\frac{2}{3}+x=\frac{1}{2}-\frac{1}{5}x\)

Logic007:

Combine \(\frac{3}{5}x\) and \(x\) u will get \(\frac{8}{5}x+\frac{2}{3}=\frac{1}{2}-\frac{1}{5}x\)

Logic007:

Add both sides by \(\frac{1}{5}x\) U will get \(\frac{9}{5}x+\frac{2}{3}=\frac{1}{2}\) Subtract both sides by \(\frac{2}{3}\) \(\frac{9}{5}x=\frac{1}{2}-\frac{2}{3}\) \(\frac{9}{5}x=-\frac{1}{6}\) \(9x(6)=-5(1)\) \(63x=-5\)

Logic007:

\(54x=-5\)*

Logic007:

Now,try to compare my working with the option given to find the correct answer